### v1

@@ Abstract
Recent advancements in text-to-image diffusion models have demonstrated their remarkable capability to generate high-quality images from textual prompts. However, increasing research indicates that these models memorize and replicate images from their training data, raising tremendous concerns about potential copyright infringement and privacy risks. In our study, we provide a novel perspective to understand this memorization phenomenon by examining its relationship with cross-attention mechanisms. We reveal that during memorization, the cross-attention tends to focus disproportionately on the embeddings of specific tokens. The diffusion model is overfitted to these token embeddings, memorizing corresponding training images. To elucidate this phenomenon, we further identify and discuss various intrinsic findings of cross-attention that contribute to memorization. Building on these insights, we introduce an innovative approach to detect and mitigate memorization in diffusion models. The advantage of our proposed method is that it will not compromise the speed of either the training or the inference processes in these models while preserving the quality of generated images. Our code is available at github.com/renjie3/MemAttn.
@@
Memorization Text-to-image Diffusion Model Cross Attention
@@ 1 Introduction
Recent advancements in diffusion models have demonstrated exceptional capabilities in image generation. Notably, text-to-image diffusion models, such as DALL-E [22] and Stable Diffusion [23], excel at producing high-quality images that adhere precisely to textual prompts. However, it is evident from increasing research that these models can memorize training data, including images and their relationship with the input textual prompts [5, 19, 18, 32]. Such a memorization issue poses a significant risk of copyright infringement of the training data and increases the possibility of privacy-sensitive information leakage [38, 11, 3].
@@
In this work, we introduce a novel perspective to understand memorization via the behavior of “cross-attention”. Cross attention has been widely used by text-to-image diffusion models, serving as the primary mechanism of selecting information from the prompts to guide diffusion generation process[23, 13, 25, 24]. Given that the memorized training images are usually triggered by the memorized textual prompts, cross attention could exhibit unique behaviors specific to memorization. Therefore, we investigate the relationship between the cross-attention and memorization. Specifically, we empirically observe a significant difference between the distributions of the cross-attention with/without memorization. One such example is demonstrated in Fig. 1. The memorized samples (Fig. 0(a) and 0(b)) tend to allocate most of the attention to the embeddings of specific tokens throughout all the diffusion steps. In contrast, non-memorization (Fig. 0(c) and 0(d)) has a more dispersed attention distribution. We deeply investigate the memorization behavior of cross attention and unveil insightful findings which pave us a way to design strategies to detect and mitigate memorization. Our detection approach is based on the quantification of the attention behavior in Sec. 4. It only requires simple calculation on attention score, and does not need repeated or additional inference operations which are often required by existing methods [5, 36]. Meanwhile, we propose an inference-time mitigation method and a training-time mitigation method that effectively reduce memorization by adjusting the attention dispersion. Similar to our detection approach, the proposed mitigation methods can preserve output quality and have little effect on image generation speed. Extensive experiments are conducted to validate our insights and the efficacy of our detection and mitigation strategies.
@@ 2 Related works
2.0.1 Memorization in generative models.
Researchers have found that generative models tend to memorize and reproduce their training data [4, 5, 2, 17, 19, 18]. Such behaviors bring significant privacy and copyright concerns [38, 11, 3]. Somepalli et al. [27, 29] show that diffusion models exhibit the behavior of producing verbatim training images. Based on this, Carlini et al. [5] develop a strategy to extract data from diffusion models by generating numerous images to assess generation density. To address these concerns, researchers have explored several strategies. Among these works, data de-duplication is recognised as an important mitigation strategy [5, 33]. Additionally, data augmentations are introduced to prevent the model overfitting the training set [7, 29]. More recently, Wen et al. [36] propose a new mitigation based on their observation that the model output of memorized prompts has a significant difference from non-memorized prompt.
@@
Cross-attention mechanisms are instrumental in most of the text-to-image diffusion models such as Stable Diffusion [23], enabling the precise generation of images that correspond to textual prompts [13, 25, 24]. Application of controlled text-to-image generation by modifying the cross-attention has been widely proposed [35, 15, 16]. For example, Hertz et al. [8] and Chefer et al. [6] have demonstrated how cross-attention can be used to refine image editing and address issues like catastrophic neglect, respectively. Besides, cross attention is also applied to text-image matching [14, 34] and explanation of text-to-image generation [30]. These works showcase the versatility and importance of cross-attention in creating semantically coherent image generation models.
@@ 3.1 Generation Process of Diffusion Models
where $\beta_{t}$ is the variance schedule. The reverse process generates new images following
@@
where $\alpha_{t}=1-\beta_{t}$, $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{t}$, $\sigma_{t}$ can be $\sqrt{\beta_{t}}$ or $\sqrt{\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}}$, and $w\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. To connect Eq. (2) with (3), $\mu_{\theta}(x_{t},t)=(x_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t))/\alpha_{t}$, and $\Sigma_{\theta}(x_{t},t)$ is determined by $\sigma_{t}$. Note that when generating new images in reverse process, we start from a random noise $x_{T}$ and end at step 00 with the output as $x_{0}$.
@@ 4 Cross Attention in Memorization
Cross attention plays an important role in text-to-image diffusion models, where it integrates the guidance of textual conditions into the image generation process. This mechanism facilitates that different segments of the generated image can pay attention to different tokens in the prompt and combine them reasonably in a global scope. In this section, we present our findings on memorization via cross attention through empirical studies. The experimental studies conducted in this section are based on SD v1.4 with the memorized prompts extracted by Webster [32] and 500 non-memorized prompts generated by ChatGPT-4 [1]. In Appendix 0.A and Sec. 6.2, we further show that the findings can also be generalized to SD v2.0.
Figure 2: Attention of beginning token
@@ 5.1 Memorization Detection
$D=\frac{1}{T_{D}}\sum_{t=0}^{T_{D}-1}E_{t}+\frac{1}{T_{D}}\sum_{t=0}^{T_{D}-1}\left|E_{t}^{\text{summary}}-E_{T}^{\text{summary}}\right|,$ (6)
where $E_{t}^{\text{summary}}=\sum_{i=N-S+1}^{N}-\bar{a}_{i}\log\left(\bar{a}_{i}\right)$ is the entropy calculated only on the summary tokens and $S$ is the number of summary tokens. The term $T_{D}$ is the number of generation steps used in calculating $D$. For SD, we choose $T_{D}=\frac{T}{5}$, which means we use the last $\frac{T}{5}$ steps of the reverse diffusion process to calculate $D$. The first term is inspired by Finding 1 that trigger tokens will cause higher entropy especially than non-memorization when $t$ is close to 00. The second term is from Finding 2 that the sum of attention score of summary tokens in memorization will reduce slower than non-memorization. In Eq. 6, we use $E_{t}^{\text{summary}}$ to replace sum to make its numerical values comparable with the first term, otherwise it becomes a negligible term. In Appendix 0.B, we show that the sum and $E_{t}^{\text{summary}}$ have consistent conclusion with Finding 2.
@@ 6 Experiment
In this section, we provide experiments to further support Finding 1 and demonstrate the effectiveness of the proposed methods. In Sec. 6.2 and Sec. 6.3, we evaluate our performance on detection and mitigation, respectively. In Sec. 6.4, we conduct ablation study on the two components of inference-time mitigation. Due to space limitation, additional experiments and details including a further verification showing the memorization’s reliance on the embeddings of trigger tokens, a validation of our findings with SD v2.0, and selection of the pre-defined threshold for training-time mitigation are elaborated in Appendix 0.A and 0.C.
@@ 6.1 Experiment Settings
Baselines. For detection, we use two baselines for comparison. The first is the method from Carlini et al. [5]. They observe that the model generates the same image for the memorized prompt when using different random seeds. Thus they propose to detect memorization by the $l_{2}$ distance of the images generated using different $n$ seeds. The second is from Wen et al. [36]. Their detection method is built upon the observation where the difference of the output between memorized prompt and empty prompt is larger than the difference between non-memorized and empty prompt. To stabilize the method, they generate the image for $n$ times. For mitigation, we compare with a the method from Wen et al. [36]. More details can be found at Sec. 0.C.1.
Evaluation metrics. The detection metrics include the area under the receiver operating characteristic curve (AUROC), the true positive rate at the false positive rate of 3% (TPR@3%FPR) and the time cost. The mitigation metrics include Similarity Score [20, 28] that measures the degree of memorization by the similarity between generated images and original training images (higher similarity means more severe memorization); Fréchet Inception Distance (FID) [9] that measures the generation quality in realism and diversity (lower FID indicates better quality); CLIP score [21] that measures fine-tuning performance (higher CLIP score means better fine-tuning performance); and the time cost. Details can be found at Sec. 0.C.2.
@@ 6.3 Memorization Mitigation
For inference-time mitigation, Fig. 5(a) presents the results of similarity score and FID with and without mitigation. The FID is calculated between generated images (including both memorized prompts from [32] and 500 non-memorized prompts) and a subset of LAION [26] with 10,000 images. Similarity Score is calculated between the generated images by memorized prompts and the training images paired with the memorized prompts. When there is no mitigation, Similarity Score is 0.7. Our method can significantly reduce Similarity Scores from 0.7 to the range of 0.25 to 0.3 (by setting $C$ from 1.1 to 1.25). Interestingly, FID is also improved after mitigation. This is likely attributable to the fact that RM and TM tend to produce memorized images that are similar to each other. Our mitigation introduces a greater diversity, thereby contributing to a reduced FID. Although the method of Wen et al. [36] has similar mitigation effect, it is slower than ours by 20% as shown in Fig. 5(b). Notably, our method maintains the same computational efficiency as performing inference without any mitigation.

### v2

@@ Abstract
Recent advancements in text-to-image (T2I) diffusion models have demonstrated their remarkable capability to generate high-quality images from textual prompts. However, increasing research indicates that these models memorize and replicate images from their training data, raising concerns about potential copyright infringement and privacy risks. In our study, we provide a novel perspective to understand this memorization phenomenon by examining its relationship with cross-attention mechanisms. We reveal that during memorization, the cross-attention tends to focus disproportionately on the embeddings of specific tokens. The diffusion model is overfitted to these token embeddings, memorizing corresponding training images. To elucidate this phenomenon, we further identify and discuss various intrinsic findings of cross-attention that contribute to memorization. Building on these insights, we introduce an innovative approach to detect and mitigate memorization in diffusion models. The advantage of our proposed method is that it will not compromise the speed of either the training or the inference processes in these models while preserving the quality of generation. Our code is available at github.com/renjie3/MemAttn.
@@
Memorization T2I Diffusion Model Cross Attention
@@ 1 Introduction
Recent advancements in diffusion models have demonstrated exceptional capabilities in image generation. Notably, text-to-image (T2I) diffusion models [25, 26] excel at producing high-quality images that adhere precisely to textual prompts. However, these models can memorize training data, including images and their relationship with the input textual prompts [7, 22, 21, 36]. Such a memorization issue poses a significant risk of copyright infringement of the training data and increases the possibility of privacy-sensitive information leakage [42, 14, 5].
@@
In this work, we introduce a novel perspective to understand memorization via the behavior of “cross-attention”. Cross attention has been widely used by text-to-image diffusion models, serving as the primary mechanism of selecting information from the prompts to guide diffusion generation process[26, 16, 28, 27]. Given that the memorized training images are usually triggered by the memorized textual prompts, cross attention could exhibit unique behaviors specific to memorization. Therefore, we investigate the relationship between the cross-attention and memorization. Specifically, we empirically observe a significant difference between the distributions of the cross-attention with/without memorization. One such example is demonstrated in Fig. 1. The memorized samples (Fig. 1(a) and 1(b)) tend to allocate most of the attention to the embeddings of specific tokens throughout all the diffusion steps. In contrast, non-memorization (Fig. 1(c) and 1(d)) has a more dispersed attention distribution. We deeply investigate the memorization behavior of cross attention and unveil insightful findings which pave us a way to design strategies to detect and mitigate memorization. Our detection approach is based on the quantification of the attention behavior in Sec. 4. It only requires simple calculation on attention score, and does not need repeated or additional inference operations which are often required by existing methods [7, 40]. Meanwhile, we propose an inference-time mitigation method and a training-time mitigation method that effectively reduce memorization by adjusting the attention dispersion. Similar to our detection approach, the proposed mitigation methods can preserve output quality and have little effect on image generation speed. Extensive experiments are conducted to validate our insights and the efficacy of our detection and mitigation strategies.
@@ 2 Related works
2.0.1 Memorization.
Researchers have found that generative models tend to memorize and reproduce their training data [6, 7, 4, 20, 22, 21], which brings significant privacy and copyright concerns [42, 14, 5]. Somepalli et al. [30, 32] show that diffusion models exhibit the behavior of producing verbatim training images. Based on this, Carlini et al. [7] develop a strategy to extract data from diffusion models by generating numerous images to assess generation density. To address these concerns, researchers have explored several strategies, such as data de-duplication [7, 37] and data augmentations [10, 32]. More recently, Wen et al. [40] propose a new mitigation based on their observation that the model output of memorized prompts has a significant difference from non-memorized prompt.
@@
Cross-attention mechanisms are instrumental in the T2I diffusion models [26], enabling the precise generation that correspond to textual prompts [16, 28, 27]. Controlled T2I generation by modifying the cross attention has been widely proposed [39, 18, 19]. For example, cross-attention can be used to refine image editing [11] and address issues like catastrophic neglect [8]. Besides, cross attention is also applied to text-image matching [17, 38] and explanation of T2I generation [33]. These works show the versatility and importance of cross attention in creating semantically coherent generation models.
@@ 3.1 Generation Process of Diffusion Models
where $\beta_{t}$ is the variance schedule. The reverse process generates images by
@@
where $\alpha_{t}=1-\beta_{t}$, $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{t}$, $\sigma_{t}$ can be $\sqrt{\beta_{t}}$ or $\sqrt{\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}}$, and $w\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. To connect Eq. (2) with (3), $\mu_{\theta}(x_{t},t)=(x_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t))/\sqrt{\alpha_{t}}$, and $\Sigma_{\theta}(x_{t},t)$ is determined by $\sigma_{t}$. Note that when generating new images in reverse process, we start from a random noise $x_{T}$ and end at step 00 with the output as $x_{0}$.
@@ 4 Cross Attention in Memorization
Cross attention plays an important role in text-to-image diffusion models, where it integrates the guidance of textual conditions into the image generation process. This mechanism facilitates that different segments of the generated image can pay attention to different tokens in the prompt and combine them reasonably in a global scope. In this section, we present our findings on memorization via cross attention through empirical studies. The experimental studies conducted in this section are based on SD v1.4 with the memorized prompts extracted by Webster [36] and 500 non-memorized prompts generated by ChatGPT-4 [3]. In Appendix A and Sec. 6.2, we further show that the findings can also be generalized to SD v2.0, transformer-based diffusion model [9], T5-based diffusion models [1, 9] and fine-tuned SD [2, 35].
@@ 5.1 Memorization Detection
$D=\frac{1}{T_{D}}\sum_{t=0}^{T_{D}-1}E_{t}-\frac{1}{T_{D}}\sum_{t=0}^{T_{D}-1}\left|E_{t}^{\text{summary}}-E_{T}^{\text{summary}}\right|,$ (6)
where $E_{t}^{\text{summary}}=\sum_{i=N-S+1}^{N}-\bar{a}_{i}\log\left(\bar{a}_{i}\right)$ is the entropy calculated only on the summary tokens and $S$ is the number of summary tokens. The term $T_{D}$ is the number of generation steps used in calculating $D$. For SD, we choose $T_{D}=\frac{T}{5}$, which means we use the last $\frac{T}{5}$ steps of the reverse diffusion process to calculate $D$. The first term is inspired by Finding 1 that trigger tokens will cause higher entropy especially than non-memorization when $t$ is close to 00. The second term is from Finding 2 that the sum of attention score of summary tokens in memorization will reduce slower than non-memorization. In Eq. 6, we use $E_{t}^{\text{summary}}$ to replace sum to make its numerical values comparable with the first term, otherwise it becomes a negligible term. In Appendix B, we show that the sum and $E_{t}^{\text{summary}}$ have consistent conclusion with Finding 2.
@@ 6 Experiment
In this section, we provide experiments to further support Finding 1 and demonstrate the effectiveness of the proposed methods. In Sec. 6.2 and Sec. 6.3, we evaluate our performance on detection and mitigation, respectively. In Sec. 6.4, we conduct ablation study on the two components of inference-time mitigation. Due to space limitation, additional experiments and details including a further verification showing the memorization’s reliance on the embeddings of trigger tokens, a validation of our findings with SD v2.0, and selection of the pre-defined threshold for training-time mitigation are elaborated in Appendix A and C.
@@ 6.1 Experiment Settings
Baselines. For detection, we use two baselines for comparison. The first is the method from Carlini et al. [7]. They observe that the model generates the same image for the memorized prompt when using different random seeds. Thus they propose to detect memorization by the $l_{2}$ distance of the images generated using different $n$ seeds. The second is from Wen et al. [40]. Their detection method is built upon the observation where the difference of the output between memorized prompt and empty prompt is larger than the difference between non-memorized and empty prompt. To stabilize the method, they generate the image for $n$ times. For mitigation, we compare with a the method from Wen et al. [40]. More details can be found at Appendix C.1.
Evaluation metrics. The detection metrics include the area under the receiver operating characteristic curve (AUROC), the true positive rate at the false positive rate of 3% (TPR@3%FPR) and the time cost. The mitigation metrics include Similarity Score [23, 31] that measures the degree of memorization by the similarity between generated images and original training images (higher similarity means more severe memorization); Fréchet Inception Distance (FID) [12] that measures the generation quality in realism and diversity (lower FID indicates better quality); CLIP score [24] that measures fine-tuning performance (higher CLIP score means better fine-tuning performance); and the time cost. Details can be found at Appendix C.2.
@@ 6.3 Memorization Mitigation
For inference-time mitigation, Fig. 6(a) presents the results of similarity score and FID with and without mitigation. The FID is calculated between generated images (including both memorized prompts from [36] and 10,000 non-memorized prompts) and a subset of LAION [29] with 10,000 images. Similarity Score is calculated between the generated images by memorized prompts and the training images paired with the memorized prompts. When there is no mitigation, Similarity Score is 0.7. Our method can significantly reduce Similarity Scores from 0.7 to the range of 0.25 to 0.3 (by setting $C$ from 1.1 to 1.25). FID has almost no change after mitigation, which means the mitigation will not influence generation quality. Although the method of Wen et al. [40] has similar mitigation effect, it is slower than ours by 20% as shown in Fig. 6(b). Notably, our method maintains the same computational efficiency as performing inference without any mitigation.
@@ 7 Conclusion
Acknowledgement
Jie Ren, Han Xu, Yaxin Li, Shenglai Zeng, and Jiliang Tang are supported by the National Science Foundation (NSF) under grant numbers CNS 2246050, IIS1845081, IIS2212032, IIS2212144, IOS2107215, DUE 2234015, DRL 2025244 and IOS2035472, the Army Research Office (ARO) under grant number W911NF-21-1-0198, the Home Depot, Cisco Systems Inc, Amazon Faculty Award, Johnson&Johnson, JP Morgan Faculty Award and SNAP.
@@ 0.A.1 Further Verification of Memorization’s Reliance on the Embeddings of Trigger Tokens
0.A.3 Extension to T5-based and Transformer-based Diffusion Models
In this subsection, we will provide results and analysis for T5-based models [9, 1], transformer-based models [9], and SD fine-tuned for high quality [2, 35]. We conduct experiments to verify that our detection and mitigation methods also work on these models, as in Tab. 2 and Tab. 3.
Table 2: Extension of detection to other T2I models
Different Architecture High-Quality Fine-tuning
  IF [1]  Pixart-$\alpha$ [9]  epiCRealism [2]  Diffusion-DPO [35]
AUROC 0.929 0.903 0.905 0.958
Table 3: Extension of mitigation to other T2I models
 w/ mitigation  w/o mitigation
epiCRealism 0.24 0.59
Diffusion-DPO 0.30 0.68
To summarize the experiments, for detection in Tab. 2 (following settings in Tab. 1 in the main paper), our method can achieve AUROC higher than 0.9 for all new models. For mitigation in Tab. 3 (following settings in Fig. 6(a) in the main paper), the inference-time mitigation reduces the similarity significantly to 0.24 and 0.30 on [2] and [35].
To explain the behavior of the models in detail, while the high-level intuition of concentrated attention on trigger tokens in memorization holds, there are some slight differences among different models. For T5-based, T5 encoder is bi-directional, which is different from the causal CLIP encoder in Stable Diffusion (SD). This means that end and padding tokens do not necessarily summarize the whole sentence. Besides, T5 encoder has no beginning tokens, while SD concentrates attention on the CLIP beginning token when focusing on denoising. Thus, Pixart-$\alpha$ [9], which is T5- and transformer-based, concentrates on the end token (This is possibly because end token contains less information and Pixart-$\alpha$ knows end token). The trigger token will distract attention from end token, which leads to disperse distribution and a high entropy. When finally using the entropy to detect memorization, we can still observe the discrepancy with/without memorization. The detection performance on Pixart-$\alpha$ shows that our main findings still holds on T5-based and transformer-based diffusion models.
For ImageFloyd (IF) [1], it is special because the cross attention is built between the image and the concatenation of prompt tokens and the image itself. The attention is in a shape of $D_{I}\times(D_{t}+D_{I})$, where $D_{I}$ is image dimension, and $D_{t}$ is prompt dimension. In the steps focusing on denoising, the image distributes most of attention to the image itself. (The image covers most of attention, while each dimension of image has a small attention.) Thus, for memorized prompts, attention will concentrate on trigger tokens with no competitor token. This leads to a low entropy in contrast to all other models.
For fine-tuned SD for high-quality, the aforementioned [2] and [35] are fined-tuned from checkpoint of SD. We find that they also memorize most of prompts of SD.